OpenStudy (anonymous):

Simplify (x – 4)(x2 + 3x + 2)

OpenStudy (johnweldon1993):

Have you heard of F-O-I-L? You take your first term in the first parenthesis...and multiply it by every term in the second parenthesis.....then...you take the second term in your first parenthesis and multiply it again by every term in the second parenthesis...so x * x² = ...? x * 3x = ...? x * 2 = ...? -4 * x² = ...? -4 * 3x = ...? -4 * 2 = ...?